### v1

@@
Phase field dislocation dynamics study of grain boundary-dislocation interactions.
@@
A new phase field dislocation dynamics formulation is presented, which couples micromechanical solvers and the time-dependent Ginzburg-Landau equation. Grain boundary (GB)-dislocation interactions are studied by describing GBs as inclusions. Grain boundary properties are computed from Molecular Statics simulations and an additional contribution to the total energy that takes into account the GB energy is considered in the calculations. Interaction of a screw dislocation with minimum energy and metastable states of low and high angle 110 symmetric tilt grain boundaries are studied. We show good agreement between predictions from our phase field dislocation dynamics formulation and molecular dynamics simulations of grain boundary-dislocation interactions.
@@ 1 Introduction
††LA-UR-25-26159New manufacturing processes such as Cold Spray [Murgas et al.,, 2024], Accumulative Roll Bonding (Severe Plastic Deformation) [Zhang et al.,, 2023], and Laser Powder Bed Fusion [Gray III et al.,, 2017] produce microstructures with complex and dense grain boundary (GB) networks, resulting in material response that varies considerably when compared to microstructures fabricated with classical manufacturing methods. During deformation, dislocations nucleate, annihilate and move through the grains encountering obstacles that affect their motion. As GBs are inherent in polycrystalline materials, they are a common impediment to dislocation motion and play an important role in controlling the material properties. The Hall-Petch relationship, which predicts that the flow stress is inversely proportional to the square root of the grain size, is the most common link between grain size (or GB density) and a mechanical property [Hall,, 1951, Petch,, 1953]. In other words, microstructures with small grains have more GBs that impede dislocation motion resulting in an increase of the yield stress. Conversely, microstructures with larger grains have lower yield stress.
@@
However, experimental observations [Nieto-Valeiras et al.,, 2024] and molecular dynamics (MD) simulations show that certain GBs do not behave as strong obstacles to dislcocation motion [Dang et al.,, 2024]. GB-dislocation interactions vary, and dislocations can be absorbed, pinned, fully or partially transmitted when impinging on a GB [Lim and Raj,, 1985, Sutton,, 1995, Zaefferer et al.,, 2003, Priester,, 2012]. The type of interaction depends on the type of dislocation, applied thermomechanical load and nature of the GB, i.e., GB misorientation, GB structure, and GB plane. Hence, by predicting the GB-dislocation interaction and engineering ‘favorable’ GBs, it is possible to enhance and tailor the in-service properties of the materials, especially during the application of new manufacturing processes where GB networks are dense [Gray III et al.,, 2017]. Current criteria for dislocation transmission across GBs are based on experimental observations and consider geometric parameters, stress state and energetics, or combinations of these, in their functional forms. Geometric criteria are based on the alignment of the slip direction, plane and intersection line between the slip-plane normal and GB plane between incoming and outgoing slip systems. Examples of geometric criteria include: [Livingston and Chalmers,, 1957], [Shen et al.,, 1986], [Luster and Morris,, 1995], residual Burgers vector [Marcinkowski and Tseng,, 1970, Bollmann,, 2012], misorientation or disorientation [Aust and Chen,, 1954, Clark and Chalmers,, 1954] and [Werner and Prantl,, 1990]. Stress-based criteria consider the Schmid factor of the incoming and outgoing slip systems [Reid,, 2016], the generalized Schimd factor [Bieler et al.,, 2014] or the resolved shear stress of the incoming and outgoing slip systems [Lee et al.,, 1989]. Combinations of geometric parameters with the accumulated shear or the Schmid factor [Bieler et al.,, 2014] have been proposed. Lastly, other criteria consider the energy change of the interaction using a line tension model [Koning et al.,, 2002]. In this model, the energy of the incoming, outgoing and residual dislocations is consider, as well as GB dislocations and GB energy barriers [Tsuru et al.,, 2009, Sangid et al.,, 2011, Li et al.,, 2020, 2022].
@@
Lower scale simulations have provided new insights into GB-dislocations interactions. MD simulations are ideal to study unit processes dictating GB-dislocation interactions over a wide range of GB structures and types of dislocation [Spearot and Sangid,, 2014, Adams et al.,, 2019, Dang et al.,, 2024, 2025, Suresh et al.,, 2023]. In Spearot and Sangid, [2014] and Adams et al., [2019], geometric criteria, residual Burgers vector and GB disorientation angle have been identified as indicators of dislocation transmission as observed experimentally; however, these studies were limited to minimum energy GB configurations. In Dang et al., [2024, 2025], GB-interactions were studied for stable and metastable GB configurations, with different reactions for stable and metastable GB configurations. These results indicate that GB structure, including minimum energy and metastable states, should be considered in the development of multiscale material models.
@@
Due to time and length scale limitations of atomistic simulations, mesoscale models are necessary to upscale information and new mechanisms to the continuum scale. Examples of mesoscale models are Discrete Dislocation Dynamics (DDD) [Li et al.,, 2009], Coupled Atomistic/Discrete Dislocation (CADD) [Dewald and Curtin,, 2006, 2007, 2011, Tadmor et al.,, 1996, Shilkrot et al.,, 2002], and Concurrent Atomistic-Continuum (CAC) [Xu et al.,, 2016, Su et al.,, 2023, Peng et al.,, 2022, Chen and Lee,, 2005]. In DDD, dislocations are represented by segments and the precision of simulations highly depends on the rules defined a priori [Cai and Bulatov,, 2004, Bertin et al.,, 2024]. In Li et al., [2009], GBs were described as geometric interfaces and the interaction was simplified to permeable or impermeable. In Bamney et al., [2021], the authors presented a DDD framework in which the structure of stable and metastable GBs were represented using dislocation arrays. However, a transmission rule needs to be defined and, e.g., in Bamney et al., [2021], the authors used the LRB criterion to model transmission [Shen et al.,, 1986]. Moreover, DDD does not represent the stacking fault of dislocations which can impact the GB-dislocation interaction, i.e., partial transmission and extended stacking faults [Dang et al.,, 2024]. CADD depends on the definition of defects in the continuum domain and the treatment of the internal boundary force that connects the quasi-continuum and continuum defect problems without the introduction of ghost forces. CADD simulations with GBs have been limited to 2D studies, hence limiting their application to special GBs with specific symmetries [Dewald and Curtin,, 2006, 2007, 2011]. CAC has been applied to study GB-dislocaiton interactions in FCC metals [Xu et al.,, 2016, Su et al.,, 2023] and dislocation pileup in a bicrystal system with square and hexagonal lattice [Peng et al.,, 2022]. CAC simulations are faster than MD simulations, however, as in other coarse-graining methodology (CAC and CADD), the time step of the simulation is imposed by the events on the atomic scale domain. Despite the technical advantages of CAC over CADD, the studies have been limited to 2D or 2.5D domains. On the other hand, CADD and CAC need to reconstruct the GB structure at the atomic scale and are limited to materials with available interatomic potentials.
@@
Another mesoscopic tool available to study GB-dislocation interactions is the phase field dislocation dynamics (PFDD) model [Koslowski et al.,, 2002, Beyerlein and Hunter,, 2016]. In this model, dislocations evolve by minimizing the total energy, hence, their interactions and reactions follow an energetically favorable pathway. Lei et al., [2013] extended the PFDD formulation to heterogeneous materials using virtual strains that describe the effect of heterogeneities and an additional evolution equation to evolve the virtual strains [Wang et al.,, 2002]. In Lei et al., [2013], the authors modeled heterogeneous materials with evolution of dislocations and voids. In Zeng et al., [2016], PFDD was used for the first time to model bimetal interfaces using PFDD with virtual strains. In Zeng et al., [2016], a misfit strain was introduced to describe the GB with a decay function that reduces the effect of the misfit strain inside the grains, while the misfit strain was derived using linear elasticity and plane stress [Hoagland et al.,, 2002], it can also be computed from molecular statics (MS) [Kharouji et al.,, 2024]. More recently, [Ma et al.,, 2022] studied an incoherent twin boundary (ITB) consisting of an array of three partial dislocations. The ITB material parameters and structure were informed by MS. GB-dislocation interaction events were in good agreement between PFDD and MD. This methodology allows the method to describe GB structures in detail and study GB-dislocation interactions precisely; however, is limited to certain systems where the GB structure can be defined in terms of well-defined dislocation arrays.
@@
The goal of this paper is twofold, first, we present a new PFDD formulation for heterogeneous materials coupled to a micromechanical solver. The results presented here use the basic scheme introduced by Moulinec and Suquet, [1994], but can be used with other schemes. The second goal is focused on the simulation of GB-dislocations interactions using an approach that can account for various GB structures without being limited to well-defined GB structures where the dislocation arrays can be determined a priori. To achieve this, we have performed simulations of bicrystals where the GB is defined as an inclusion. The simulations parameters are informed by MS [Mishra et al.,, 2024], and compared to recent MD simulations [Dang et al.,, 2025]. In particular, we focus on the observations made by Dang et al., [2025]: GB-dislocation interaction varies depending on the applied stress, misorientation, and whether it is a minimum energy or metastable GB. MD simulations show that interactions can vary and involve only the leading partial during transmission (partial transmission). To capture this, PFDD simulations were performed using partial dislocations [Hunter et al.,, 2013]. In Section 2, the PFDD formulation is presented. This section describes how the micromechanical problem is solved, and how the total strain is used in the elastic energy terms that describe the evolution of the dislocations. Next, Section 3 presents the PFDD simulation results of dislocation reactions with different GBs in Cu. Finally, concluding remarks are given in Section 4.
@@ 2 Phase field dislocation dynamics formulation coupled with the Lippmann-Schwinger equation
In the PFDD formulation, dislocations are described using a scalar order parameter, , that is defined on each active slip system, . and correspond to a perfect crystal and a crystal that has undergone slip, while non-integer values represent a distorted crystal structure (i.e., the location of the dislocation core). The coupled formulation uses two different sets of equations, one to obtain the total strain through solving the Lippmann-Schwinger equation and the other to obtain the evolution of dislocations in the system following the pseudo-algorithm shown in Figure 1. The different components of the algorithm are described in sections below.
@@ 2.1 Dislocation mechanics solving the Lippmann-Schwinger equation
where the total strain is split into a fluctuation term and a prescribed volume average term corresponding to the homogeneous strain, and is the plastic strain. The plastic strain depends on the order parameter as [Koslowski et al.,, 2002, Beyerlein and Hunter,, 2016]
where is the Burgers vector magnitude, is the interplanar spacing, is the number of active slip systems, and, and are the slip direction and slip plane normal of slip system . Introducing a reference material with stiffness and using the decomposition of the elastic strain, the stress may be redefined as
@@
Introducing the Green’s function, the above equation can be solved in the form of an integral Lippmann-Schwinger equation for the unknown strain:
where denotes spatial convolution and is the modified Green tensor defined as the second derivative of the Green’s function . In this work, Equation 8 is solved using the basic scheme proposed by Moulinec and Suquet, [1994] and reference stiffness tensor , where and are the minimum and maximum elastic properties, see Algorithm 1 in the Supplementary Material.
@@ 2.2 Dislocation evolution using the time-dependent Ginzburg-Landau (TDGL) kinetic equation
The total energy is composed of three contributions [Koslowski et al.,, 2002, Beyerlein and Hunter,, 2016]:
where , and are the elastic, lattice and external energy. Note that in some PFDD formulations there is an additional gradient energy term representing the excess energy generated by the non-uniformity of the dislocation [Xu et al.,, 2019, Shen and Wang,, 2003, Kim et al.,, 2021]. The elastic energy accounts for the elastic interactions between dislocations and is defined as
To define the elastic energy of the heterogeneous system, the elastic strain can be derived from Equation 4 using Green’s functions and the decomposition of the total strain:
and in Fourier space
where is the wave number vector and the total strain is obtained by solving the Lippmann-Schwinger equation, see Equation 8. By applying the Plancherel theorem and using the definition of the elastic strain in Fourier space, the elastic energy is defined as
where the interaction matrices , and are defined as
The lattice energy represents the energy required to move the dislocation core through the crystal lattice by the breaking and reforming of atomic bonds. It can be expressed using different forms such as piecewise quadratic functions [Koslowski et al.,, 2002], sinusoidal functions [Xu et al.,, 2020, Peng et al.,, 2020] or Fourier series [Hunter et al.,, 2011, Schoeck,, 2001, Shen and Wang,, 2004], depending on the simulation parameters:
where are the periodic potentials. For simulations in this work, it is necessary to use partial dislocations. For this, we consider the slip plane and three different slip directions with order parameters, , and , that correspond to the directions , and , respectively [Hunter et al.,, 2013].
@@
where is the applied stress. The homogeneous strain is determined by the boundary conditions. For a stress-free system, the homogeneous strain is determined by minimizing the total strain energy with respect to the homogeneous strain
@@
where , , , and . In a homogeneous system, Equation 20 is redefined as which simplifies to the definition of the external energy defined in the homogeneous PFDD formulation [Jin and Khachaturyan,, 2001].
The system evolves using a time-dependent Ginzburg-Landau (TDGL) kinetic equation, which minimizes the total energy
where is a coefficient that represents the mobility of the dislocations.
The lattice energy in Equation 17 is redefined as
where is implemented as a Fourier series [Hunter et al.,, 2011, Schoeck,, 2001, Shen and Wang,, 2004]:
The expansion coefficients and can be obtained by fitting the -surface computed using MD or DFT, in this work we used the same coefficients computed by Ma et al., [2022].

### v2

@@
Phase field dislocation dynamics formulation coupled with Fourier based micromechanics solver and its application to grain boundary-dislocation interactions.
@@
A new phase field dislocation dynamics formulation for homogeneous and heterogeneous materials is presented, which couples micromechanical solvers and the time-dependent Ginzburg-Landau equation. The strain fields are obtained from the micromechanical solver by solving the Lippmann-Schwinger equation, and then used to define energy terms to model the evolution of the dislocations. Grain boundary (GB)-dislocation interactions are studied using the coupled PFDD formulation and by describing GBs as inclusions. GB energy and stiffness tensors are computed from molecular statics simulations and a newly proposed lattice energy term that is dependent on the GB energy is considered in the calculations. Interaction of a screw dislocation with minimum energy and metastable states of low and high angle 110 symmetric tilt grain boundaries are studied. We show good agreement between predictions from our phase field dislocation dynamics formulation and molecular dynamics simulations of grain boundary-dislocation interactions.
@@ 1 Introduction
††LA-UR-25-26159New manufacturing processes such as Laser Powder Bed Fusion [Gray III et al., 2017, Rodríguez-Barber and Pérez-Prado, 2025], Accumulative Roll Bonding (Severe Plastic Deformation) [Zhang et al., 2023], and Cold Spray [Murgas et al., 2024] produce microstructures with complex and dense grain boundary (GB) networks, resulting in material response that varies considerably when compared to microstructures fabricated with classical manufacturing methods and the possibility to tailor properties by controlling the GBs. During deformation, dislocations nucleate, annihilate and move through the grains encountering obstacles that affect their motion. As GBs are inherent in polycrystalline materials, they are a common impediment to dislocation motion and play an important role in controlling the material properties. The Hall-Petch relationship, which predicts that the flow stress is inversely proportional to the square root of the grain size, is the most common link between grain size (or GB density) and a mechanical property [Hall, 1951, Petch, 1953].
However, experimental observations [Nieto-Valeiras et al., 2024] and molecular dynamics (MD) simulations show that certain GBs do not behave as strong obstacles to dislocation motion [Dang et al., 2024]. Rather, GB-dislocation interactions vary, and dislocations can be absorbed, pinned, fully or partially transmitted when impinging on a GB [Lim and Raj, 1985, Sutton, 1995, Zaefferer et al., 2003, Priester, 2012]. The type of interaction depends on the type of dislocation, applied thermo-mechanical load and nature of the GB, including the misorientation of the neighboring grains and (atomic or misfit) structure of the GB. Hence, by predicting the GB-dislocation interaction and engineering ‘favorable’ GBs, it is possible to enhance and tailor the in-service properties of the materials, especially during the application of new manufacturing processes where GB networks are dense [Gray III et al., 2017, Murgas et al., 2024, Zhang et al., 2023, Rodríguez-Barber and Pérez-Prado, 2025].
@@
Current criteria for dislocation transmission across GBs are based on experimental observations and consider geometric parameters, stress state and energetics, or combinations of these, in their functional forms. Geometric criteria are based on the alignment of the slip direction , plane and intersection line between the slip-plane normal and GB plane normal between incoming (subscript ) and outgoing (subscript ) slip systems. Examples metrics of different geometric criteria include: [Livingston and Chalmers, 1957], [Shen et al., 1986], [Luster and Morris, 1995], residual Burgers vector [Marcinkowski and Tseng, 1970, Bollmann, 2012], misorientation or disorientation [Aust and Chen, 1954, Clark and Chalmers, 1954] and [Werner and Prantl, 1990]. Stress-based criteria consider the Schmid factor of the incoming and outgoing slip systems [Reid, 2016], the generalized Schimd factor [Bieler et al., 2014] or the resolved shear stress of the incoming and outgoing slip systems [Lee et al., 1989]. Combinations of geometric parameters with the accumulated shear or the Schmid factor [Bieler et al., 2014] have been proposed. Lastly, other criteria consider the energy change of the interaction using a line tension model [Koning et al., 2002]. In this model, the energy of the incoming, outgoing and residual dislocations is considered, as well as GB dislocations and GB energy barriers [Tsuru et al., 2009, Sangid et al., 2011, Li et al., 2020, 2022].
@@
Lower scale simulations have provided new insights into GB-dislocations interactions that motivate the need to account for the structure of the GB itself. MD simulations are ideal to study unit processes dictating GB-dislocation interactions over a wide range of GB structures and types of dislocation [Spearot and Sangid, 2014, Adams et al., 2019, Dang et al., 2024, 2025, Suresh et al., 2023]. In Spearot and Sangid [2014] and Adams et al. [2019], geometric criteria, residual Burgers vector and GB disorientation angle have been identified as indicators of dislocation transmission as observed experimentally; however, these studies were limited to minimum energy GB configurations. In Dang et al. [2024, 2025], GB-dislocation interactions were studied for stable and metastable GB configurations, and different reactions were observed for stable and metastable GB configurations. These results indicate that GB structure, including minimum energy and metastable states, should be considered in the development of multiscale material models.
@@
Due to time and length scale limitations of atomistic simulations, mesoscale models are necessary to upscale information and new mechanisms to the continuum scale. Examples of mesoscale models are Discrete Dislocation Dynamics (DDD) [Li et al., 2009], Coupled Atomistic/Discrete Dislocation (CADD) [Dewald and Curtin, 2006, 2007, 2011, Tadmor et al., 1996, Shilkrot et al., 2002], and Concurrent Atomistic-Continuum (CAC) [Xu et al., 2016, Su et al., 2023, Peng et al., 2022, Chen and Lee, 2005]. In DDD, dislocations are comprised of segments and the precision of simulations highly depends on the rules defined a priori [Cai and Bulatov, 2004, Bertin et al., 2024]. In Li et al. [2009], GBs were described as geometric interfaces and the interaction was simplified to permeable or impermeable. In Bamney et al. [2021], the authors presented a DDD framework in which the structure of stable and metastable GBs were represented using dislocation arrays. However, a transmission rule needs to be defined and, e.g., in Bamney et al. [2021], the authors used the LRB criterion to model transmission [Shen et al., 1986]. Moreover, these studies did not account for partial dislocations and the stacking fault between them, which can impact the GB-dislocation interaction [Dang et al., 2024].
CADD depends on the definition of defects in the continuum domain and the treatment of the internal boundary force that connects the quasi-continuum and continuum defect problems without the introduction of ghost forces. CADD simulations with GBs have been limited to 2D studies, hence limiting their application to special GBs with specific symmetries [Dewald and Curtin, 2006, 2007, 2011]. CAC has been applied to study GB-dislocation interactions in face centered cubic (FCC) metals [Xu et al., 2016, Su et al., 2023] and dislocation pileup in a bicrystal system with square and hexagonal lattices [Peng et al., 2022]. CAC simulations are faster than MD simulations, however, as in other coarse-graining methodology (CAC and CADD), the time step of the simulation is imposed by the events on the atomic scale domain. Despite the technical advantages of CAC over CADD, the studies have been limited to 2D or 2.5D domains. On the other hand, CADD and CAC need to reconstruct the GB structure at the atomic scale and are limited to materials with available interatomic potentials.
@@
Another mesoscopic tool available to study GB-dislocation interactions is the phase field dislocation dynamics (PFDD) model [Koslowski et al., 2002, Beyerlein and Hunter, 2016]. In this model, dislocations evolve through minimization of the total system energy, hence, their interactions and reactions follow an energetically favorable pathway. Lei et al. [2013] extended the PFDD formulation to heterogeneous materials using a virtual strain formulation to describe the effect of heterogeneities within the total system energy [Wang et al., 2002]. This type of PFDD formulation uses an Eshelby inclusion approach to account for the heterogeneous elastic constants within the system. Thus, the entire system is given the elastic constants of the matrix material and the regions that are meant to be inhomogeneous have a non-zero virtual strain (also sometimes called eigen strain) field to account for the differences. In other words, there are non-zero values of the virtual/eigen strains only in the inclusions, and as the dislocation evolves the nine components of the virtual/eigen strain evolve every time step using an additional kinetic equation. In Lei et al. [2013], the evolution of dislocations in a heterogeneous material with voids were considered. In Zeng et al. [2016], PFDD was used for the first time to model bimetal interfaces also using the virtual strain formulation. In addition to employing the virtual strains to account for elastic inhomogeneities, a misfit strain was also introduced to describe the lattice mismatch at the bimetal interface. This strain component was derived using linear elasticity and plane stress [Hoagland et al., 2002]. Such a misfit strain could also be computed from molecular statics (MS) [Kharouji et al., 2024] as well as disclination mechanics using the fast Fourier transform (FFT) [Berbenni and Taupin, 2018] or the finite element method [Zhang et al., 2018], and use it as an additional eigen strain in the PFDD model. More recently, Ma et al. [2022] studied an incoherent twin boundary (ITB) consisting of an array of three partial dislocations. The ITB material parameters and structure were informed by MS. GB-dislocation interaction events were in good agreement between PFDD and MD. This methodology allows the method to describe GB structures in detail and study GB-dislocation interactions precisely; however, is limited to certain systems where the GB structure can be defined in terms of well-defined dislocation arrays.
@@
Clearly, FFT-based models have been able to successfully model a diverse set of problems [Moulinec and Suquet, 1994, Lebensohn, 2001, Bertin et al., 2015, Berbenni et al., 2014, Michel et al., 2001, Chen et al., 2019, 2015, Vidyasagar et al., 2017] and carries advantages such as a computational complexity of compared to of the finite element method, reduced memory allocation needs (matrix free formulation), and the possibility to use images as inputs as pixels (voxels) are used in 2D (3D). Thus, we first present a new PFDD formulation for heterogeneous materials coupled to a FFT-based micromechanical solver. The coupled formulation we present is consistent with the homogeneous PFDD formulation [Koslowski et al., 2002]; however, compared to the virtual strain formulation [Lei et al., 2013], it does not require virtual strain fields to evolve the strain within the inclusions. The results presented here use the non-linear FFT micromechanical solver known as the basic scheme introduced by Moulinec and Suquet [1994] with an eigen strain that represents the plastic deformation from the dislocation. However, different solvers could be used if desired. Note that the basic scheme requires a lot of iterations to converge if the ratio of the elastic properties tend to infinity or zero. Second, we present simulations of GB-dislocation interactions using an approach that can account for general GB structures, removing the limitation to well-defined GB structures where the structure must be determined a priori. This approach is demonstrated through simulations of bicrystals where the GB is defined as an inclusion. The simulations parameters are informed by MS [Mishra et al., 2024], and compared to recent MD simulations [Dang et al., 2025]. In particular, we focus on the observations made by Dang et al. [2025]: GB-dislocation interaction varies depending on the applied stress, misorientation, and whether it is a minimum energy or metastable GB. MD simulations show that interactions can vary and involve only the leading partial during transmission (partial transmission). To capture this, PFDD simulations were performed using partial dislocations [Hunter et al., 2013]. In Section 2, the PFDD formulation is presented. This section describes how the micromechanical problem is solved, and how the total strain is used in the elastic energy terms that describe the evolution of the dislocations. Next, Section 3 presents the PFDD simulation results of dislocation reactions with different GBs in Cu. Finally, concluding remarks are given in Section 4.
@@ 2 Phase field dislocation dynamics formulation coupled with the Lippmann-Schwinger equation
In the PFDD formulation, dislocations are described using a scalar order parameter, , that is defined on each active slip system, . and correspond to a perfect crystal and a crystal that has undergone slip, respectively, while non-integer values between 0 and 1 represent a distorted crystal structure (i.e., the location of the dislocation core). This coupled formulation uses two different sets of equations, one to obtain the total strain through solving the Lippmann-Schwinger (LS) equation and the other to obtain the evolution of dislocations in the system following the pseudo-algorithm shown in Figure 1. Note that every time step the staggered scheme follows this sequence: (i) the mechanical problem is solved, (ii) the energy terms are defined, and (iii) the dislocation evolves. The three different components of the algorithm are described in the subsections below.
@@ 2.1 Dislocation mechanics solving the Lippmann-Schwinger equation
where the total strain is split into a fluctuation term and a prescribed volume average term corresponding to the homogeneous strain, and is the plastic strain. The plastic strain depends on the order parameter as [Koslowski et al., 2002, Beyerlein and Hunter, 2016, Nabarro, 1951]
where is the Burgers vector magnitude, is the interplanar spacing, is the number of active slip systems, and, and are the slip direction and slip plane normal of slip system . The sum over means that the plastic strain is proportional to the number of active slip systems. In this formulation, all of the order parameters are aligned with perfect dislocation slip systems in FCC metals, as this is the material system studied later in the Results Section (Section 3). Partial dislocations are modeled through linear combinations of the 3 perfect dislocation order parameters on a given slip plane. We note that the PFDD method is not limited to FCC systems, but has also been applied to body centered cubic (BCC) [Fey et al., 2024, 2022], hexagonal close packed (HCP) [Albrecht et al., 2021], and semiconductor systems [Kim et al., 2022]. For the case of HCP materials, the Burgers vector magnitude and interplanar distance vary with the active slip systems, and would require that and both also be a function of the slip system, . However, this is not the case for the FCC system considered here, and these superscripts are dropped for convenience.
Introducing a reference material with stiffness and using the decomposition of the elastic strain, the stress may be redefined as
@@
Introducing the Green’s function, the above equation can be solved in the form of an integral Lippmann-Schwinger equation for the unknown strain [Kröner, 1971]:
where denotes spatial convolution and is the modified Green’s tensor defined as the second derivative of the Green’s function and if . In this work, Equation 8 is solved using the basic scheme proposed by Moulinec and Suquet [1994] and reference stiffness tensor , where and are the minimum and maximum elastic properties, see Algorithm 1 in the Supplementary Material.
@@ 2.2 Dislocation evolution using the time-dependent Ginzburg-Landau (TDGL) kinetic equation
The total energy is composed of three contributions [Koslowski et al., 2002, Beyerlein and Hunter, 2016]:
where , and are the elastic, lattice and external energy. Note that in some PFDD formulations there is an additional gradient energy term representing the excess energy generated by the non-uniformity of the dislocation [Xu et al., 2019, Shen and Wang, 2003, Kim et al., 2021], which is neglected in the current work. The elastic energy accounts for the elastic interactions between dislocations and is defined as
with being the volume of the simulation domain. To define the elastic energy of the heterogeneous system, the elastic strain can be derived from the decomposition of the total strain in Equation 2 and the definition of the total strain in Equation 8. Instead of applying the Plancherel theorem [Plancherel and Leffler, 1910] to obtain the elastic energy in the Fourier space, in this work we compute the elastic energy in the real space using Equation 10.
The lattice energy represents the energy required to move the dislocation core through the crystal lattice by the breaking and reforming of atomic bonds. It can be expressed using different forms such as piecewise quadratic functions [Koslowski et al., 2002], sinusoidal functions [Xu et al., 2020, Peng et al., 2020] or Fourier series [Hunter et al., 2011, Schoeck, 2001, Shen and Wang, 2004], depending on the simulation parameters:
where are the periodic potentials. For simulations in this work, it is necessary to account for partial dislocations. For this, we consider the slip plane and three different slip directions with order parameters, , and , that correspond to the directions , and , respectively [Hunter et al., 2013].
The lattice energy in Equation 11 can then be redefined as
where is implemented as a Fourier series [Hunter et al., 2011, Schoeck, 2001, Shen and Wang, 2004]:
The expansion coefficients and can be obtained by fitting the -surface computed using MD or DFT; in this work we used the same coefficients computed by Ma et al. [2022].
@@
where is the applied stress. The homogeneous strain is determined by the boundary conditions. For a system in equilibrium, the homogeneous strain is determined by minimizing the total strain energy with respect to the homogeneous strain
@@
where , , , and . For an applied strain ,the external energy is defined as . In a homogeneous system, Equation 16 is redefined as which simplifies to the definition of the external energy defined in the homogeneous PFDD formulation [Jin and Khachaturyan, 2001, Koslowski et al., 2002].
Finally, the dislocation order parameter evolves using a time-dependent Ginzburg-Landau (TDGL) kinetic equation, which minimizes the total energy
where is a coefficient that represents the mobility of the dislocations, is the energy density and a variational derivative is used. The variational derivative can be transformed into a partial derivative as [Giaquinta and Hildebrandt, 2004] (c.f. Chapter 1, Section 2.1)
as the gradient term is neglected in this work, the TDGL equation is
This equation is discretized in time following the forward Euler method as
where the subscripts and refers to the current and next time step. More details about the formulation are shown in Supplementary Section 2.